OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)?

OpenStudy (anonymous):

(x − 2)2 + (y + 3)2 = 5 (x + 2)2 + (y − 3)2 = 25 (x − 2)2 + (y + 3)2 = 25 (x − 2)2 − (y + 3)2 = 5

OpenStudy (anonymous):

it is c

OpenStudy (anonymous):

there are two ways to solve it 1 simply put the points (-2, 0)in the eq

OpenStudy (anonymous):

2 simply find the distance between them that is radius and use the circle formula

OpenStudy (anonymous):